OpenStudy (anonymous):

2x - 5y = 12 whats the sloope?

OpenStudy (anonymous):

\[\frac{2}{5}\]

OpenStudy (anonymous):

solve for y and it will be displayed for you \[2x-5y=12\] \[5y=2x-12\] \[y=\frac{2}{5}x -\frac{12}{5}\]

OpenStudy (anonymous):

y=mx+n then m is slope